Judy bought a coat at a 20% discount sale for $72. What was the origional price of the coat?
Found 2 solutions by user_dude2008, mathgranny:
Answer by user_dude2008(1862)

72/0.8 = 90 $

Answer by mathgranny(13)

Judy bought a coat at a 20% discount sale for $72. What was the original price of the coat?
Imagine yourself buying a coat.
The items mentioned are 20% discount sale , original price .
Since we don't use o for a variable (looks like 0)
let's use r for regular price(or original Price)
When something is on sale we subtract something from he regular price to get the sale price.

Fill in the blank: regular price - ______ = sale price.
Regular price - actual discount amt. = sale price
r - d = s
We need to find regular price. That means we are either given the sale price and the discount amt. or we have the info to find them.

There are 2 numbers in the problem.
What does $72 represent. We are told sale price = $72.
Now how do we use 20% to find discount amt.
Discount amt is the small amt taken away from the reg price to get sale price. d = 20% of r so d = .20r


reg. price - discount = sale price
r - d = s
r - .20r = 72
1r - .20r = 72
.80r = 72
I see multiplication , so I divide
r = 72/.80 = $90
Ask yourself if this answer is reasonable.
Is a 20% discount large or small. I'd say small.
If reg price is $90 and sale price is $72 is this a small decrease or a large one. I'd say small.